OpenStudy (anonymous):

in 1991 the life expectancy of males in a certain country was 64.6 years . in 1995 it was 68.4 years. let E represent the life expectancy in year T represent the number of years since 1991 the liner function E(t) that fits data is e(t)=__ t+__ round to the nearest 10th use the function to predict the life expectancy of males in 2005 E(14)= round to the nearest 10th

OpenStudy (bahrom7893):

I think I did this problem before.......

OpenStudy (anonymous):

hmm

OpenStudy (bahrom7893):

x=1 y= 64.6 x=5 y=68.4

OpenStudy (anonymous):

e(t)=__ t+__?

OpenStudy (anonymous):

E(14)=

OpenStudy (bahrom7893):

y-y1 = M(x-x1) M = (x-x1)/(y-y1) M = (5-1)/(68.4-64.6) M = 4 / 3.8 = 1.0526

OpenStudy (bahrom7893):

y - 64.6 = 1.0526(x-1) or E(t) = 1.0526x-1.0526+64.6 E(t) = 1.0526x+63.5474

OpenStudy (bahrom7893):

wait round to the tenth so: E(t) = 1.1x+63.5

OpenStudy (bahrom7893):

*sorry meant E(t) = 1.1t+63.5 E(14) = 1.1*14+63.5

OpenStudy (bahrom7893):

E(14) = 78.9

OpenStudy (bahrom7893):

that's the final answer.

OpenStudy (anonymous):

thank you sooooooooooo much!!!!